OpenStudy (anonymous):

If a and b are two sides of a right triangle, find the length of the hypotenuse, c. a=6, b=2, c=?

OpenStudy (mertsj):

a^2+b^2=c^2

OpenStudy (anonymous):

well using the pythagorean theorm; a^2 + b^2 = c^2 so pluggin it in 6^2 + 2^2 = 40 sqrt c^2 = 2 sqrt 10 do you get it? ^.^

OpenStudy (anonymous):

ohhhh! i get it now! :) thanks! you're the person that's helped me more than anyone and not just gaven me the answer, you helped me understand the answer! thank you so so much :) @SovaySable

OpenStudy (anonymous):

You're very welcome! ^.^

OpenStudy (anonymous):

@SovaySable can you help me with some more please? :)))

OpenStudy (anonymous):

lol sure!

OpenStudy (anonymous):

okay! simplify:

OpenStudy (anonymous):

okay square roots are basically getting rid of repeats in any given equation; so let's take the first one: sqrt 9x^3 we know tht 9 is a repeat of 3 so square root that x^3 has a repeat in it as well because x^2 times x equals x^3 therefore we break down the x^3 and square root the x squared outta there leaving just x under the sqrt doing this for the rest of the variables we get: 3xsqrtx + 10xsqrtx - 6xsqrtx and since they are all like variables we just do the simple math across ^.^ and wala! your answer will be 7x sqrt x ^.^ I never thought math could be fun >.> but do you get it?? any questions?

OpenStudy (anonymous):

yay i get it! :)))) i'm having sort of trouble with some other ones too. but thanks for explaining it to me!

OpenStudy (anonymous):

do you get this one? it's another simplifying problem

OpenStudy (anonymous):

@SovaySable

OpenStudy (anonymous):

NO ACTUALLY it's not simplifying! solve for x!

OpenStudy (anonymous):

hmm...did you do imaginary numbers? cuz I think this equation has an imaginary number in it

OpenStudy (anonymous):

@SovaySable can you simplify this please?

OpenStudy (anonymous):

sure its x^3 sqrt x =)

OpenStudy (anonymous):

@SovaySable i need your help if you can help me! :(

OpenStudy (anonymous):

haha sure; one more though okay cuz i have to sleep.

OpenStudy (anonymous):

divide & simplify!

OpenStudy (anonymous):

and: The distance between the points (1,8) and (4,4) is equal to 5. True False

OpenStudy (anonymous):

okay for the first one which is a radical we can see that both the denominator and numerator is square rooted so thus we combine both under one. so now we have: sqrt 2x^3/x^5 now do basic algebric division and you get x^-2 or 2/x^2 which is the simplest form it'd make more sense if you write it out =) any questions??

OpenStudy (anonymous):

and for the last question; true; 5 is the distance between points! Hope this helps! =)